Unemployment rate: natural rate of unemployment-0.5 x (output gap) 1. Use Okuns Law to estimate the following: a) The current US unemployment rate is around 10%, and level of inflation is close to 0%. If the natural rate of unemployment is 6%, what can we say about the output gap (or at what production capacity is the economy is operating relative to the potential capacity is the économy is operating relative to the potential output?)? The current US unemployment rate is around 13%, and the GDP Deflator falls by 4%. If the production capacity is currently at 85% of its potential output, what is the approximate natural rate of unemployment? b)
Answer #1

a) Unemployment rate = natural rate of unemployment - 0.5*(output gap)

or, 10% = 6%-0.5*(Output gap) or, 05.*(output gap) = 4% or output gap = 8%

b) Current unemployment rate = 13% GDP deflator or inflation rate further falls by 4%. Production capacity is 85% of potential output which means output gap is 15%

therefore, 13% = Natural rate of unemployment - 0.5*15% or, 13% = Natural rate of unemployment - 7.5%

or, Natural rate of unemployment = (13-7.5)%= 5.5%.